OpenStudy (janu16):

Wendell is looking over some data regarding the strength, measured in Pascals (Pa), of some building materials and how the strength relates to the length. The data are represented by the exponential function f(x) = 2^x, where x is the length. Explain how he can convert this equation to a logarithmic function when strength is 8 Pascals.

OpenStudy (phi):

it sounds like they want you to write 2^x = 8 then take the log of both sides

OpenStudy (janu16):

so log 2^x= log 8? @phi

OpenStudy (phi):

yes. they may want you to simplify that to x= log(8)/log(2) though that does not look like a function to me.

OpenStudy (janu16):

ohok. what do i do after that?

OpenStudy (janu16):

@phi

OpenStudy (phi):

they *might* want \[ \log_2(x) = 3 \] but I'm guessing.

OpenStudy (janu16):

but i have to show work before i write 3 as the answer so how would i show work?

OpenStudy (phi):

if you use log base 2 (rather than base 10) \[\log_2 2^x= \log_2 8\] and knowing 8 is 2^3 you an say \[ x \log_2 2 = \log_2 2^3 \\ x = 3 \] but I am not sure that is what they are asking for.

OpenStudy (janu16):

they probbaly are. thanks so much!

OpenStudy (janu16):

so log2 8=log2 2^3 is in logarithmic form right?